Three coins are tossed in the air. what is the probability that exactly two of them turn up heads?

Answer by stanbon(26297)

Three coins are tossed in the air. what is the probability that exactly two of them turn up heads?
------------------
P(x=2) = 3C2(1/2)^2*(1/2) = 3*(1/4)(1/2) = 3/8
==================================================
Cheers,
Stan H.